Time value  - An Iowa state savings bond can be converted to ​$250 at maturity 5 years from purchase. If the state bonds are to be competitive with U.S. savings​ bonds, which pay 3​% annual interest​ (compounded annually), at what price must the state sell its​ bonds? Assume no cash payments on savings bonds prior to redemption. Ignore taxes.

Question: The state must sell its bonds for ​$? ​(Round to the nearest​ cent.)
Answer #1

Single Payment Present Worth Factor (PWF)=(P/F i, N)=1/((1+i)^N)

i=Interest Rate=3%=0.03

N=Number of years=5

PWF=1/(1.03^5)=0.862609

Current Selling Price =$250*PWF=$250*0.862609=$215.65

The state must sell its bonds for

$215.65